# A Position Paper on the Automatic Generation of Machine Learning Leaderboards

### v1

@@ 8 Conclusion
In the position paper, we provide the first overview of ALG research, which reveals substantial diversity in problem framing and benchmarking practices. To address this fragmentation, we propose an ALG unified conceptual framework and present ALG benchmarking guidelines. Furthermore, our first overview of ALG research to date revealed that the scope of current leaderboards is limited. Therefore, one key recommendation in our list of challenges and new directions for ALG is to expand leaderboard coverage. Future leaderboards should report all results, including baselines, ablations, and method variations, and enrich tuples with broader metadata (e.g. hyperparameters) to create a more informative resource. In support of this initiative, a continually updated reading list is maintained in a GitHub repository181818https://github.com/RoelTim/ML-leaderboard-position-paper.
@@
A limitation of this paper is the scope, as we solely focus on the automatic generation of ML leaderboards. We note that other disciplines also report experimental outcomes, although the nature of the experimental procedures may differ. For example, ghosh2024toward explores finetuning LLMs for schema-based information extraction in material science. Another example is wang2024scidasynth, which introduced SciDaSynth, an interactive system using LLMs to extract and synthesise structured knowledge from the scientific literature in the form of tables.
@@
This research is subject to the governance by the ethics board of ANONYMOUS. We note that our proposal for AI research is to facilitate decision-making by users, as opposed to complete automation of tasks. We note that data mining activities for scientific literature should comply with the terms and conditions of the publishers disseminating published work, noting that scientific text mining is often consider to be fair use of copyright material. The use of AI for data mining in this case is on public domain material.
Appendix A Related Work Beyond ALG
@@
Entity and relation extraction from scientific papers gained attention in 2017 with the SemEval-2017 ScienceIE task, which focused on identifying key elements like processes, tasks, and materials in publications augenstein2017semeval. The SemEval-2018 Task 7 advanced this by classifying relationships such as “uses”, “compares”, and “improves” between scientific concepts buscaldi2018semeval. Datasets like SciERC luan2018multi, TDMSci Hou2021TDMSci, and Dmdd pan2023dmdd further support entity extraction research. The methods developed for scientific entity and relationship extraction can be leveraged to generate scientific leaderboards automatically.
@@
A scientific leaderboard compares methods, highlighting the best-performing one. It is a specific case of structured scientific information comparison and meta-analysis. Research has focused on extracting structured information without emphasizing leaderboards. For example, ghosh2024toward explored LLMs for schema-based information extraction in material science. walker2023extracting improved extraction of experimental procedures using fine-tuned language models, while wang2024scidasynth introduced SciDaSynth, an interactive system using LLMs to extract and synthesise structured knowledge from scientific literature.
Appendix B Problem Framing Details
@@
| Methodology | Document Representation |
| TDMS-IE Hou2019Leaderboards | DocTAET*, SC |
| ORKG-TDM Kabongo2021Mining | DocTAET |
| ORKG-LB Kabongo2023ORKG | DocTAET |
| PI Graph singh2019automated | Citation Tables |
| AXCELL Kardas2020AXCELL | Full Paper & Tables |
| SciREX-IE Jain2020SCIREX | Full Paper |
| TELIN Yang2022TELIN | Full Paper & Tables |
| TDMS-PR kabongo2024effective | DocREC† |
| MS-PR Singh2024LEGOBENCH | Full Paper |
| TDMR-PR sahinuc2024efficient | Full Paper & Tables |
| * Hou2019Leaderboards perform ablation studies with variations of DocTAET. | |
| † kabongo2024effective compare the performance of three document representations: DocREC, DocTAET, and the Full Paper. | |
Appendix C Methodology Details
@@
Hou2019Leaderboards propose TDMS-IE, a methodology to automatically extract task, dataset, metric, score tuples from research papers. The first step of TDMS-IE is extracting the document representation and the score context from the research paper. The document representation, DocTAET, covers the title, abstract, experimental setup, and table information. The title and abstract help predict the task, while the experimental setup and table information assist in identifying the dataset and metric. A second document-based structure, the score context, SC, represents contents from tables, since the work relies on table-based (and formatting, i.e., bold font) heuristics to generate candidate tuples. The SC captures the table caption and column headers corresponding to each bold-faced numeric score in each table of the research paper. This is used in conjunction with formatting-based heuristics to identify candidates for the best score of a task, dataset, metric tuple.191919For example, bold-faced scores are most likely to be best score. Hou2019Leaderboards frame the problem as a natural language inference (NLI) task using two entailment models: 1) DocTAET-TDM and 2) SC-DM. Each model generates a tuple hypothesis (a Task-Dataset-Metric, or TDM, tuple for DocTAET-TDM; a Score-Dataset-Metric tuple for SC-DM), by searching for candidate argument combinations from a “taxonomy” (that is, a knowledge base) of previously observed tuples. A fine-tuned BERT model (for NLI) predicts whether a candidate tuple can be inferred from DocTAET, inferring links between the paper’s text and the predefined canonical labels for the Task, Dataset, and Metric, as represented in the taxonomy. For instance, the model can recognise that "Rg-2" and "ROUGE-2" refer to the same metric. Similarly, the SC-DM infers entailment relationships between the SC document representations and dataset-metric tuples. Both models use the BERT model limited to 512 tokens devlin2019bert, although newer models with larger token capacities may improve performance.
@@
Appendix D Dataset Details
@@
| Entities | Format | Annotations* | Unk. | Train Stats. | Test Stats. | Val. Stats. | ||||||||||||
| Paper | V | T | D | M | S | Md | LaTeX | HA | PwC | NLPP | Ann. | #P | #T | #P | #T | #P | #T | |
| ORKG-PwC Dataset | ||||||||||||||||||
| Kabongo2021Mining | v1 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | 2,831† | 11,724† | 1,228† | 5,060† | - | - |
| Kabongo2021Mining | v2 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | 3,753† | 11,724† | 1,608† | 5,060† | - | - |
| Kabongo2023ORKG | v3 | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | 587† | 9,614† | 270† | 4,096† | - | - |
| Kabongo2023ORKG | v4 | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 2,946† | 9,614† | 1,262† | 4,096† | - | - |
| Kabongo2023ORKG | v5 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | 587† | 9,614† | 270† | 4,096† | - | - |
| Kabongo2023ORKG | v6 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | 2,946† | 9,614† | 1,262† | 4,096† | - | - |
| Kabongo2023ZeroShot | v7# | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | - | - | 1,000 | 1,925 | - | - |
| NLP-TDMS Dataset | ||||||||||||||||||
| Hou2019Leaderboards | v1 | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | 124 | 325 | 118 | 281 | - | - |
| Hou2019Leaderboards | v2 | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | 170 | 325 | 162 | 281 | - | - |
| Kardas2020AXCELL | v3 | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✓ | 170 | 325 | 162 | 281 | - | - |
| PwC-LB Dataset | ||||||||||||||||||
| Kardas2020AXCELL | v1 | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | ‡ | ‡ | 516 | 2,802 | ‡ | ‡ |
| Yang2022TELIN | v2 | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | - | - | 516 | 2,802 | - | - |
| SciREX Dataset | ||||||||||||||||||
| Jain2020SCIREX | ✓ | ✓ | ✓ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ | 438 | 438 | 438 | ||||||
| TDMS-Ctx Dataset | ||||||||||||||||||
| kabongo2024effective | v1§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 11,807 | 402,409 | 1,326 | 33,863 | - | - |
| kabongo2024effective | v2§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 12,388 | 415,788 | 1,401 | 34,799 | - | - |
| kabongo2024effective | v3§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 10,058 | 415,788 | 1,105 | 31,213 | - | - |
| kabongo2024effective | v4§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 11,807 | 402,409 | 746 | 14,604 | - | - |
| kabongo2024effective | v5§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 12,388 | 415,788 | 789 | 14,800 | - | - |
| kabongo2024effective | v6§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 10,058 | 415,788 | 595 | 14,273 | - | - |
| LEGOBench Dataset | ||||||||||||||||||
| Singh2024LEGOBENCH | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | - | - | 43,105 | - | - | ||
| SciLead Dataset | ||||||||||||||||||
| sahinuc2024efficient | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | - | - | 43 | - | - | ||
@@
| * For annotations, we distinguish between human annotations (HA), Papers with Code (PwC) and NLP Progress (NLPP), however PwC includes partial human annotation, and domain experts fully curated NLP Progress via GitHub pull requests. Use LaTeX if available; otherwise, default to PDF. ‡Different data for training (unlabelled arXiv papers and segmented tables) and validation (linked results). †Two-fold cross-validation: 70% train, 30% test, with averaged results. 9,847 leaderboards, and the number of papers is unspecified. § v1–v3 are few-shot experiment datasets with document representations: v1 (DocFULL), v2 (DocREC), and v3 (DocTAET). v4–v6 are zero-shot experiment datasets with the same representations: v4 (DocFULL), v5 (DocREC), and v6 (DocTAET). # the same data source as v2, but with updated timestamps and no overlap with v2. An average of 5 tuples annotations per paper. Unspecified, with 138 unique tuples reported. | ||||||||||||||||||
Appendix E Definitions of Metrics

### v2

@@ 8 Conclusion
In the position paper, we provide the first overview of ALG research, which reveals substantial diversity in problem framing and benchmarking practices. To address this fragmentation, we propose an ALG unified conceptual framework and present ALG benchmarking guidelines. Furthermore, our first overview of ALG research to date revealed that the scope of current leaderboards is limited. Therefore, one key recommendation in our list of challenges and new directions for ALG is to expand leaderboard coverage. Future leaderboards should report all results, including baselines, ablations, and method variations, and enrich tuples with broader metadata (e.g. hyperparameters) to create a more informative resource. In support of this initiative, a continually updated reading list is maintained in a GitHub repository.191919https://github.com/RoelTim/ML-leaderboard-position-paper
@@
A limitation of this paper is the scope, as we solely focus on the automatic generation of ML leaderboards. We note that other disciplines also report experimental outcomes, although the nature of the experimental procedures may differ. For example, ghosh2024toward explore finetuning LLMs for schema-based information extraction in material science. Another example is wang2024scidasynth, which introduce SciDaSynth, an interactive system using LLMs to extract and synthesise structured knowledge from the scientific literature in the form of tables.
While this position paper does not include new experiments, it aims to establish the foundational scaffolding required for rigorous future evaluations. We therefore provide clear evaluation setups and guidelines that, for future research, can be directly applied to assess existing and future ALG systems.
@@
This research is subject to the governance by the ethics board of the Commonwealth Scientific and Industrial Research Organisation (CSIRO). We note that our proposal for AI research is to facilitate decision-making by users, as opposed to complete automation of tasks.
Appendix A Acronyms
| Acronym | Full form |
| ALG | Automatic Leaderboard Generation |
| ML | Machine Learning |
| NLP | Natural Language Processing |
| LLM | Large Language Model |
| EMNLP | Conference on Empirical Methods in Natural Language Processing |
| PTE | Predefined Typed Entities |
| PTT | Predefined Typed Tuples |
| DocTAET | Document representation: Title, Abstract, Experimental Setup, Tables |
| DocREC | Document representation: Results, Experiments, Conclusion |
| ORKG | Open Research Knowledge Graph |
| PwC | Papers with Code |
| SciREX | Scientific Research Information Extraction (dataset) |
| LEGOBench | Leaderboard Generation Benchmark |
| SciLead | Scientific Leaderboard (dataset) |
| NLI | Natural Language Inference |
| F1 | Harmonic mean of Precision and Recall |
| LR | Leaderboard Recall |
| PC | Paper Coverage |
| RC | Result Coverage |
| AO | Average Overlap |
| LaTeX | Document preparation system |
| OCR | Optical Character Recognition |
Appendix B Related Work Beyond ALG
@@
Entity and relation extraction from scientific papers gained attention in 2017 with the SemEval-2017 ScienceIE task, which focused on identifying key elements like processes, tasks, and materials in publications augenstein2017semeval. The SemEval-2018 Task 7 advanced this by classifying relationships such as “uses”, “compares”, and “improves” between scientific concepts buscaldi2018semeval. mondal-etal-2021-end built a knowledge graph from NLP papers by extracting four types of relations: “evaluatedOn” (associating tasks with datasets), “evaluatedBy” (associating tasks with evaluation metrics), as well as “coreferent” and “related” relations, which capture connections among entities of the same type. Datasets like SciERC luan2018multi, TDMSci Hou2021TDMSci, and Dmdd pan2023dmdd further support entity extraction research. The methods developed for scientific entity and relationship extraction can be leveraged to generate scientific leaderboards automatically.
@@
A scientific leaderboard compares methods, highlighting the best-performing one. It is a specific case of structured scientific information comparison and meta-analysis. Research has focused on extracting structured information without emphasising leaderboards. For example, pramanick-etal-2023-diachronic leveraged a causal discovery algorithm to identify the TDMM (task, dataset, metric, method) entities associated with a specific task and assess their causal influence on the task’s research trends. Numerical quantity extraction has also been explored in scientific text for summarisation purposes 10.1145/3539618.3591808.
Work related to AI methods for publications from other scientific disciplines may also influence ALG. ghosh2024toward explored LLMs for schema-based information extraction in material science. walker2023extracting improved the extraction of experimental procedures using fine-tuned language models. Other prior work has demonstrated the ability to extract methods and processes in scientific text for life sciences Wan2010SupportingSummariser; Wan2009WhettingContext and material science yang-2022-piekm.
wang2024scidasynth introduced SciDaSynth, an interactive system using LLMs to extract and synthesise structured knowledge from scientific literature. Recently, pramanick-etal-2025-nature proposed a method for automatically extracting, categorizing, and quantitatively analyzing contribution statements in research papers. Knowledge graph generation has been explored in other disciplines, like astronomy 10.1007/978-3-031-35320-8_15; timmer2023KG.
Appendix C Problem Framing Details
@@
| Methodology |
| TDMS-IE Hou2019Leaderboards | DocTAET*, SC | ||
| ORKG-TDM Kabongo2021Mining | DocTAET | ||
| ORKG-LB Kabongo2023ORKG | DocTAET | ||
| PI Graph singh2019automated | Citation Tables | ||
| AxCell Kardas2020AXCELL | Full Paper & Tables | ||
| SciREX-IE Jain2020SCIREX | Full Paper | ||
| TELIN Yang2022TELIN | Full Paper & Tables | ||
| TDMS-PR kabongo2024effective | DocREC† | ||
| MS-PR Singh2024LEGOBENCH | Full Paper | ||
| TDMR-PR sahinuc2024efficient | Full Paper & Tables | ||
| * Hou2019Leaderboards perform ablation studies with variations of DocTAET. | |||
| † kabongo2024effective compare the performance of three document representations: DocREC, DocTAET, and the Full Paper. | |||
Appendix D Methodology Details
@@
Hou2019Leaderboards propose TDMS-IE, a methodology to automatically extract task, dataset, metric, score tuples from research papers. The first step of TDMS-IE is extracting the document representation and the score context from the research paper. The document representation, DocTAET, covers the title, abstract, experimental setup, and table information. The title and abstract help predict the task, while the experimental setup and table information assist in identifying the dataset and metric. A second document-based structure, the score context, SC, represents contents from tables, since the work relies on table-based (and formatting, i.e., bold font) heuristics to generate candidate tuples. The SC captures the table caption and column headers corresponding to each bold-faced numeric score in each table of the research paper. This is used in conjunction with formatting-based heuristics to identify candidates for the best score of a task, dataset, metric tuple.202020For example, bold-faced scores are most likely to be best score. Hou2019Leaderboards frame the problem as a natural language inference (NLI) task using two entailment models: 1) DocTAET-TDM and 2) SC-DM. Each model generates a tuple hypothesis (a Task-Dataset-Metric, or TDM, tuple for DocTAET-TDM; a Score-Dataset-Metric tuple for SC-DM), by searching for candidate argument combinations from a “taxonomy” (that is, a knowledge base) of previously observed tuples. A fine-tuned BERT model (for NLI) predicts whether a candidate tuple can be inferred from DocTAET, inferring links between the paper’s text and the predefined canonical labels for the Task, Dataset, and Metric, as represented in the taxonomy. For instance, the model can recognise that "Rg-2" and "ROUGE-2" refer to the same metric. Similarly, the SC-DM infers entailment relationships between the SC document representations and dataset-metric tuples. Both models use the BERT model limited to 512 tokens devlin2019bert, although newer models with larger token capacities may improve performance.
@@
Appendix E Dataset Details
@@
| Entities | Format | Annotations* | Unk. | Train Stats. | Test Stats. | Val. Stats. | ||||||||||||||
| Paper |
T | D | M | S | Md | LaTeX | HA | PwC | NLPP |
#P | #T | #P | #T | #P | #T | |||
| ORKG-PwC Dataset | ||||||||||||||||||||
| Kabongo2021Mining | v1 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | 2,831† | 11,724† | 1,228† | 5,060† | - | - | ||
| Kabongo2021Mining | v2 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | 3,753† | 11,724† | 1,608† | 5,060† | - | - | ||
| Kabongo2023ORKG | v3 | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | 587† | 9,614† | 270† | 4,096† | - | - | ||
| Kabongo2023ORKG | v4 | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 2,946† | 9,614† | 1,262† | 4,096† | - | - | ||
| Kabongo2023ORKG | v5 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | 587† | 9,614† | 270† | 4,096† | - | - | ||
| Kabongo2023ORKG | v6 | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | 2,946† | 9,614† | 1,262† | 4,096† | - | - | ||
| Kabongo2023ZeroShot | v7# | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | - | - | 1,000 | 1,925 | - | - | ||
| NLP-TDMS Dataset | ||||||||||||||||||||
| Hou2019Leaderboards | v1 | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✓ | ✗ | 124 | 325 | 118 | 281 | - | - | ||
| Hou2019Leaderboards | v2 | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | 170 | 325 | 162 | 281 | - | - | ||
| Kardas2020AXCELL | v3 | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✓ | 170 | 325 | 162 | 281 | - | - | ||
| PwC-LB Dataset | ||||||||||||||||||||
| Kardas2020AXCELL | v1 | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | ‡ | ‡ | 516 | 2,802 | ‡ | ‡ | ||
| Yang2022TELIN | v2 | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | - | - | 516 | 2,802 | - | - | ||
| SciREX Dataset | ||||||||||||||||||||
| Jain2020SCIREX | ✓ | ✓ | ✓ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ | 438 | 438 | 438 | ||||||||
| TDMS-Ctx Dataset | ||||||||||||||||||||
| kabongo2024effective | v1§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 11,807 | 402,409 | 1,326 | 33,863 | - | - | ||
| kabongo2024effective | v2§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 12,388 | 415,788 | 1,401 | 34,799 | - | - | ||
| kabongo2024effective | v3§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 10,058 | 415,788 | 1,105 | 31,213 | - | - | ||
| kabongo2024effective | v4§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 11,807 | 402,409 | 746 | 14,604 | - | - | ||
| kabongo2024effective | v5§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 12,388 | 415,788 | 789 | 14,800 | - | - | ||
| kabongo2024effective | v6§ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | 10,058 | 415,788 | 595 | 14,273 | - | - | ||
| LEGOBench Dataset | ||||||||||||||||||||
| Singh2024LEGOBENCH | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | - | - | 43,105 | - | - | ||||
| SciLead Dataset | ||||||||||||||||||||
| sahinuc2024efficient | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | - | - | 43 | - | - | ||||
@@
| * For annotations, we distinguish between human annotations (HA), Papers with Code (PwC) and NLP Progress (NLPP), however PwC includes partial human annotation, and domain experts fully curated NLP Progress via GitHub pull requests. Use LaTeX if available; otherwise, default to PDF. ‡Different data for training (unlabelled arXiv papers and segmented tables) and validation (linked results). †Two-fold cross-validation: 70% train, 30% test, with averaged results. 9,847 leaderboards, and the number of papers is unspecified. § v1–v3 are few-shot experiment datasets with document representations: v1 (DocFULL), v2 (DocREC), and v3 (DocTAET). v4–v6 are zero-shot experiment datasets with the same representations: v4 (DocFULL), v5 (DocREC), and v6 (DocTAET). # the same data source as v2, but with updated timestamps and no overlap with v2. An average of 5 tuples annotations per paper. Unspecified, with 138 unique tuples reported. | ||||||||||||||||||||
Appendix F Definitions of Metrics